OpenStudy (anonymous):

You have challenged four other students to come up with expressions equal to x2. Their submissions are below. Mike Kelly Scott Kim Using complete sentences explain the steps in simplifying the expression of each student and identify who is correct.

OpenStudy (anonymous):

To simplify Kim's expression: First write out the expression: 5sqrtx*x^2/5*x^3/5*x^4/5 Next get rid of the radical: The square root of x = 1/2 but because there is a 5th root the x inside of the square root is x to the 1/5 power or x^1/5. So now we have x^1/5*x^2/5*x^3/5*x^4/5 Add the powers: x^1/5+2/5+3/5+4/5 which equals x^10/5 which simplifies to x^2 let me get the rest

OpenStudy (anonymous):

To simplify Scott's expression: First write out the expression: (6sqrtx)^3 The sqaure root of x is equal to x^1/2 but because there is a 6 instead of the given 2 6sqrtx is equal to x^1/6. Next you are going to multiply 3 into the x factor: x^1/6*3= x^6/3 which fully simplified is x^2

OpenStudy (anonymous):

.To simplify Mike's expression: First write out the expression: x^8/5*x^4/5 all over x^2/5 Multiply the top exponents together by adding the powers: x^8/5*x^4/5=x^12/5 The expression now look like this: x^12/5 all over x^2/5 To divide you subtract the powers: 12/5-2/5 which equals x^10/5 which simplifies further to x^2

OpenStudy (anonymous):

To simplify Kelly's expression: First write out the expression: 7sqrtx^2/3*x^5*x^6 You want to get the square root into exponential form before you can multiply the expressions together: 7sqrt^2/3=x^2/21. Once you have that you can multiply your powers: x^5*x^6*x^2/21 Add the powers: 5/1+6/1+2/21= 13/21 which can not be further simplified so Kelly was incorrect. Her final answer was x^13/21

OpenStudy (anonymous):

Thank you so much i been having so much trouble on this.

OpenStudy (anonymous):

your welcome